OpenStudy (anonymous):

On a backpacking trip through Europe, Troy spent 20 Swiss francs to stay one night in a youth hostel. If the foreign exchange rate between the U.S. dollar and the Swiss franc is 1:2.2, how much did Troy spend in U.S. dollars?

OpenStudy (anonymous):

1/2.2 = x/20 (now cross multiply) (2.2)(x) = (1)(20) 2.2x = 20 x = 20/2.2 x = 9.09 I believe he spent $9.09 You might want to wait till you get a second opinion....I think I did it right but I am not 100% sure.